OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5). A. 23 B. –53 C. –9 D. 3

OpenStudy (anonymous):

its nothing but f(g(-5)) g(-5) = -4*(-5) + 3 = 20+3 = 23 So we need f(g(-5)) = f(23) = -2*23 - 7 = -46-7 = -53

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

My pleasure. Having fun helping ppl out here :)